GENETICS PROBLEMS-4 (10 points total). SHOW PUNNETT SQUARES FOR EACH PROBLEM. NONE OF THESE PROBLEMS INVOLVES LINKAGE OR CROSSING OVER USE GENETIC TERMINOLOGY! 1. The ability to taste the chemical phenylthiocarbamide is an autosomal (not X linked) dominant phenotype, and the inability to taste it is recessive. Marlene is a taster whose father is a non-taster. Abdul is a taster who in a previous marriage had a non- taster daughter. Marlene is pregnant and Abdul is the father. Show this cross between Marlene and Abdul using a Punnett square. What is the probability that Marlene and Abduls child will be a non-taster girl? (HINT: this is a. b. a dihybrid problem) From the Punnett square, what are the expected phenotypes of their children? STATE NUMBERS AND TYPES c. a. b. C.
Answer #1

1. Let's represent the ability and inability to taste as T and t respectively.

Marlene is a taster, whose father was a non-taster. This means she has atleast one T allele for tasting, but as her father had both t alleles, she must have atleast one of those from her father. So her genotype must be Tt.

Abdul is a taster and has a non-tasting daughter, meaning he contributed a recessive gene. He too must be heterozygous (Tt)

a. The cross would like: Tt x Tt

T t
T TT Tt
t Tt tt

b. Probability of having a non-tasting girl:

TtXX x TtXY

TX TX tX tX
TX TTXX TTXX TtXX TtXX
TY TTXY TTXY TtXY TtXY
tX TtXX TtXX ttXX ttXX
tY TtXY TtXY ttXY ttXY

Probability of having a non-tasting girl - 2/16 = 1/8 or 12.5%

c. Expected phenotypes:

Male, tasting - 6/16 = 3/8 or 37.5%

Male, non-tasting - 2/16 = 1/8 or 12.5%

Female, tasting - 6/16 = 3/8 or 37.5%

Female, non-tasting - 2/16 = 1/8 or 12.5%